9. One hundred men and one hundred women are surveyed to determine if the percent of women that lie about their weight is the same as the percent of men, or if the percent of women is higher than the percent of men. The hypothesis test to use is ______________.

a. independent group means, population standard deviations known
b. two proportions
c. independent group means, population standard deviations unknown
d. matched or paired samples

10. Fourteen cars from two different models of a car manufacturer are compared to see if the smaller model has a higher average MPG than the larger model. From years of testing these models, it is known that the standard deviation of MPG is 3.4 MPG for the smaller model and 4.0 MPG for the larger one. From this survey, the sample means and standard deviations are 28.2 and 3.0, and 22.6 and 5.1, respectively. If we conduct a hypothesis test, the p-value is ______________.

a. 0.0004
b. 0.0002
c. 0.0000
d. 0.0010
Answer #1

(9)

Correct option:

b.   two proportions

Explanation:
The 2 independent proportions are: (i) the proportion of women that lie about their weight and (ii) the proportion of men that lie about their weigh

(10)

Correct option:

c.   0.0000

Explanation:

H0: \mu 1\leq \mu 2

HA: \mu 1> \mu 2

n1 = 14

T1 = 28.2

\sigma1 = 3.4

n2 = 14

T2 = 22.6

\sigma2= 4.0

Test Statistic is given by:

Z=\frac{28.2-22.6}{\sqrt{\frac{3.4^{2}}{14}+\frac{4.0^{2}}{14}}}=3.991

By Technology,

p - value = 0.0000